OpenStudy (boldgeorge):

Solve the following system of linear equations. 4x+3y=13 5x+y-8=0 Medal for anyone who can help me!

OpenStudy (anonymous):

Make y the subject in one of the equations and substitute it instead of y in the other equation

OpenStudy (jhannybean):

You have two equations, 1 of which you can solve for "y" and substitute in for the other. Let's take 5x+y-9=0 for example. Isolate y on the left side. \[\large 5x+y-8=0\]subtract -5x from 5x and add +8 to both sides. \[\large -5x+5x+y-8+8=0-5x+8\]\[\large y=-5x+8\] now that you have solved for y with 1 equation, we can substitute it into the other equation and solve for a variable. \[\large 4x+3(-5x+8)=13\]\[\large 4x-15x+24=13\]isolate the x's on one side. \[\large-11x = -11\]\[\large x= 1\] use this found value of x into one of two equations to solve for y and you will have both yoru variables found. :) \[\large 5(1)+y-8=0\]\[\large y-3=0\]\[\large y-3+3=0+3\]\[\large y=?\]